Suppose you have a summer reading list with 15 books and you have to read 3. How many different combinations of books can you read?
Answer by jim_thompson5910(35256)

There are 15*14*13 = 2,730 ways to read 3 books where order matters.

Since order does not matter, we divide by 3! = 3*2*1 = 6 to get 2,730/6 = 455

There are 455 different combinations.